OpenStudy (anonymous):

Janice needs to read 2/3 of her book by Friday. So far she has read 1/2 of it. What part of the book must she still read by Friday? A. 1/6 B. 5/12 C. 2/5 D. 1/3

OpenStudy (anonymous):

1/2 * 2/3 = 1/3

OpenStudy (anonymous):

D. 1/3

OpenStudy (anonymous):

k.

OpenStudy (anonymous):

Thanks!!! ^-^

OpenStudy (anonymous):

um, he's wrong the correct answer is A. 1/6 you do subtraction not multiplication, so it would be 2/3 - 1/2 which is 4/6-3/6=1/6

OpenStudy (anonymous):

I agree with Eureka70

OpenStudy (anonymous):

First Find the Common Denominator, Then Subtract, And when you do that The answer you get is 1/6

OpenStudy (anonymous):

A.1/6